OpenStudy (anonymous):

Evaluate x − 2y − 5z when x = −3, y = 1, and z = −2

OpenStudy (anonymous):

Jus tplug in and solve

OpenStudy (kirbykirby):

Just substitute the values in the original equation: -3 -2(1) -5(-2) = -3-2+10 = 5

OpenStudy (anonymous):

which will give you 5 as kirbykirby said.